Ticket Number: HEA-NOC/20070921-3 Ticket Status: OPEN
Ticket Type: unscheduled Resolver: HEA-NOC
Ticket Opened: 20070921 13:39 UTC+1 Problem Start: 20070921 12:59 UTC+1
Ticket Update: 20070921 13:39 UTC+1 Problem End: 20070921 13:15 UTC+1
12:59:54.819 tcd-sw1-vlan2 714: %STORM_CONTROL-3-FILTERED: A Broadcast
storm detected on Gi1/0/15. A packet filter action has been applied on the
Broadcast traffic would have been throttled as per configuration to protect
the network. Minimal effect to service.
20070921 13:34 GAR Checked if storm was ongoing, but it had already
finished. Will monitor to see if this recurs.
This ticket can be monitored at http://www.hea.net/tickets/20070921-3
HEAnet Limited. Registered in Ireland, No. 275301.